# Curiosity-Driven Co-Development of Action and Language in Robots Through Self-Exploration

### v2

@@ Abstract
Human infants acquire language and action co-developmentally, achieving remarkable generalization capabilities from only a minimal number of learning examples. In contrast, recent large language models require exposure to billions of training tokens to achieve such generalization. What mechanisms underlie such efficient developmental learning in humans? This study addresses this question through simulation experiments in which robots learn to perform various actions corresponding to imperative sentences (e.g., push red cube) via trials of self-guided exploration. Our approach integrates the active inference framework with reinforcement learning, enabling curiosity-driven developmental learning. The simulations yielded several nontrivial findings: i) Curiosity-driven exploration combined with motor noise substantially outperforms learning without curiosity. ii) Simpler, prerequisite-like actions emerge earlier in development, while more complex actions involving these prerequisites develop later. iii) Rote pairing of sentences and actions occurs before the emergence of compositional generalization. iv) Generalization is drastically improved as the number of compositional elements increases. These results shed light into possible mechanisms underlying efficient co-developmental learning in infants and provide computational parallels to findings in developmental psychology.
@@ Introduction
A central question in both cognitive science and artificial intelligence is how humans and artificial systems can acquire competencies for language and motor command in a co-developmental manner, despite having access to only limited learning experiences. This question is exemplified in human infants, who achieve remarkable generalization with sparse input. This is a stark contrast to large-scale models which rely on massive training corpora, to reach similar capabilities. This raises the issue of what mechanisms enable such efficient developmental learning.
@@
In linguistic terms, compositionality refers to the ability to construct novel configurations by systematically combining elements such as verbs, adjectives, and nouns. Generalization enables infants to apply learned components flexibly, allowing for the production and interpretation of utterances that have not been directly encountered previously. Although the number of possible compositions grows multiplicatively with the vocabulary size (i.e., number of verbs $\times$ number of adjectives $\times$ number of nouns), infants achieve generalization after experiencing only a small subset of learning examples. This suggests that the effective sample complexity is closer to the sum of elements rather than their product. This phenomenon is closely related to the “poverty of the stimulus” problem articulated by Chomsky (?), which asks how learners generalize so effectively given severely sparse input.
@@
To investigate these mechanisms, one promising approach is to reconstruct developmental learning processes in machines and robots. The field of developmental robotics has long pursued this line of research, aiming to replicate human-like learning trajectories in embodied systems (?, ?, ?, ?). However, relatively few studies have focused on the co-development of language and motor control under conditions of stimulus poverty. Existing work has primarily examined associative mappings between linguistic input and motor commands in one-shot or supervised batch learning schemes (?, ?, ?, ?). These approaches neglect the self-directed, developmental context of infant learning.
@@
In this study, we propose a self-exploratory learning framework of robots in which reinforcement learning is incorporated with the active inference framework (?, ?, ?), enabling curiosity-driven exploration. Our approach to integrate reinforcement learning with active inference was originally inspired by the work of Kawahara et al. (?). In our model, originally introduced in (?), motor commands are reinforced by two intrinsic rewards: curiosity (seeking unpredictable sensory consequences) and motor entropy (seeking random movements). Motor commands are also reinforced by extrinsic rewards for successfully achieving goal tasks specified by given imperative sentences. Importantly, our previous experiments in maze navigation demonstrated that the combination of curiosity and entropy is crucial for enhancing self-exploration, as agents achieved significantly improved exploratory behaviors under this dual-intrinsic reward scheme. Our approach aligns with broader research on self-exploration in machine learning, in which agents are intrinsically rewarded for taking motor commands that increase unpredictability or information gain (?, ?).
@@
A simulated mobile robot equipped with a manipulator arm, vision sensor, and distributed tactile sensors learns to generate motor movements in response to imperative sentences presented during each trial. These sentences are systematically composed of verbs, nouns, and adjectives, enabling evaluation of generalization performance under different levels of compositional complexity.
The model architecture employed in this study is based on our previous work (?) with key modifications to accommodate multi-modal sensorimotor integration. (See details in the Materials and Methods section.) Figure 1 presents the model architecture, which is composed of two main components: a forward model and an actor part.
Figure 1: The proposed model architecture. The model consists of predictor, actor, and critic.
The forward model learns to predict the next sensation $o_{t+1}$ based on the current sensation $o_{t}$ and the executed motor command $a_{t}$. The sensation includes pixel-based vision, tactile sensation, arm joint proprioception, and voice for the sentences. To address the hidden state problem and probabilistic nature of the environment, the prediction is performed contextually and stochastically using the random latent variables $z_{t}$ and the deterministic $h^{q}_{t}$. The random latent variables were allocated separately for each sensory modality, while the deterministic variables were shared. The actor module generates the next motor command $a_{t+1}$ based on the deterministic latent variable $h^{q}_{t}$, which integrates current sensation $o_{t}$.
The overall flow is: with an imperative sentence given by the tutor, the robot attempts to achieve the specified goal by generating a sequence of motor commands. Meanwhile, the forward model predicts the next sensation by inferring the posterior probability distribution $q(z_{t}|o_{t},h_{t-1})$ of the random latent variable $z$. This inference is conducted by minimizing the evidence free energy $F$ (Eq. 1). This consists of the complexity term represented by Kullback–Leibler divergence (KLD) between the estimated posterior and the prior, and the accuracy term as shown in the free energy principle (FEP). (See more details on the evidence free energy and expected free energy in the “Free Energy Principle, Active Inference, and Kawahara Model” section of the Supplementary Materials.)
$F_{\psi,t}=\underbrace{D_{KL}[q(z_{t}|o_{t},h_{t-1})||p(z_{t}|h_{t-1})]}_{\text{Complexity}}-\underbrace{\mathbb{E}_{q(z_{t})}[\log p(o_{t+1}|h_{t})]}_{\text{Accuracy}}.$ (1)
The forward model is trained incrementally by optimizing the learning parameters $\psi$ in the direction of minimizing the evidence free energy. The actor generates motor command sequences in the direction of minimizing the expected free energy $G$ (Eq. 2). This consists of the complexity term, extrinsic reward term, and the entropy term through reinforcement learning of motor commands.
$G(a_{t})$ $=-\underbrace{D_{KL}[q(z_{t}|o_{t},h_{t-1})||p(z_{t}|h_{t-1})]}_{\text{complexity}}-\underbrace{r(s_{t},a_{t})}_{\text{Extrinsic Reward}}-\underbrace{\mathcal{H}(\pi_{\phi}(a_{t}|h_{t-1}))}_{\text{Entropy}}$ (2)
It is interesting to note that minimizing evidence free energy $F$ minimizes the complexity term while minimizing expected free energy $G$ maximize the same complexity term. This means that motor commands are generated in the direction of maximizing the information gain represented by KLD between the estimated posterior and the prior after the motor command execution. This generates curiosity-driven exploration wherein the agent seeks out previously unencountered sensorimotor experiences. On the other hand, the learning parameters in the forward model is updated in the direction of minimizing the same complexity term, representing the latent conflict that is generated by novel experiences encountered during curiosity-driven exploration. Therefore, both processes of self-exploration and the forward model learning are racing each other.
@@
This study tested the following hypotheses through simulation experiments: H1: Curiosity combined with motor entropy enhances the performance of developmental learning. H2: Primitive actions are acquired earlier, followed by more complex, prerequisite-dependent actions. H3: In the early phase, actions are generated only for exactly learned sentences, but in later phases, the system generalizes to novel, unlearned compositions. H4: Generalization performance improves as the scale of compositionality in the task increases.
@@ Results
We created a robot like a truck crane in a physics simulator along with a set of objects with 5 different shapes each of which can be with 6 different colors (see Figure 2).
Figure 2: The simulated robot and a set of objects to act on. Top: the robot has two wheels and an arm with two joints. The design is similar to a truck crane. Bottom, left to right: a red pillar, a green pole, a blue dumbbell, a cyan cone, and a magenta hourglass. The color yellow is not pictured here.
@@
For each trial episode, a task goal was given in terms of an imperative sentence composed with verb, adjective, and object noun. Possible words used for them are shown in Table 1.
@@
Verb Adjective Object noun
@@
Table 1: The English words used for imperative sentences specifying goals.
@@
At each step, the robot receives visual sensation, proprioception for the arm, tactile sensation , and two types of voices: the command voice and the tutor-feedback voice. The command voice takes the format of the imperative sentence described previously, and it comes every step continuously from the beginning. On the other hand, the feedback voice arrives whenever the robot achieves one of possible goals even if the achieved goal is not the one told by the command voice, and it informs which goal has been achieved actually in the same format with the command voice. This potentially enhances the forward model to learn about own motor motions. Finally, when the goal specified by the command voice is achieved, a reward is provided. Each trial episode ran for 30 steps, or terminated when the specified goal is achieved.
@@
The experiment examined effects of different levels of curiosity to the developmental learning processes using the basic setup. In the basic setup, full compositions of words (Table ) were used to generate the imperative sentences. However, the training was conducted using only 60 imperative sentences (33 percentage) out of 180 possible sentences. 120 untrained sentences were used for generalization test. For ten agents with different random seeds, the whole developmental learning process was iterated for 60000 epochs. The generalization test with unlearned imperative sentences was conducted for every 50 epochs.
@@
Figure 3 shows the development of the generalization test performances in terms of success rate for goals specified by unlearned imperative sentences which are plotted for different action categories with different levels of curiosity. Shades areas represent 99% confidence intervals.
@@
Figure 3: Rolling success-rates for unlearned goals for agents with different levels of curiosity.
@@
The plots shows that the performance was improved significantly as the curiosity level was increased. Especially, the case of all actions with the all curiosity level shows the average success rate for unlearned goals reached a quite high value of 90 percentage even though the learning was conducted only for 33 percentage of all possible compositions. (See a video: https://www.youtube.com/watch?v=ipjXg53f-zk) showing examples of the robot behaviors with the all curiosity level in the intermediate phase and the final phase in the course of development. It can be also seen that some action categories developed faster than other action categories. Especially in the all curiosity case, “watch” developed the fastest, “be near” did as the second, “push forward” as the third, and “push left,” “push right” and “touch the top” developed much later. This implies that simpler prerequisite-type actions develop earlier, and more complex actions requiring those prerequisite actions develop later. Actually, an action of watching an object should be prerequisite for all other object-targeted actions including an action of moving near by an object, which should be prerequisite again for actions of directly manipulating an object like pushing left/right an object or touching the top of it. Our observation accords with this.
@@
Next, Figure 4 (A) shows the success rate comparison between learned and unlearned goals under the all curiosity condition for each action category. These plots show that the test performance for learned goals developed significantly faster than the case for the unlearned goals. This indicates that actions are generated only for exactly learned goal compositions in the early phase, but the system generalizes to novel, unlearned composition ones in the later phase.
@@
The experimental results are shown in Figure 4. Shades areas represent 99% confidence intervals.
Figure 4: Rolling success-rates for agents with different numbers of tasks, colors, and shapes.
@@
It can be seen that although the learned goal test cases show equally high performance for all scales of compositionality, the generalization test for unlearned goal case shows significantly lower performance as the compositionality scale decreases. This indicates that the generalization performance severely depends on the scale of compositionality in learning examples.
Further analysis on the experiment results
Some analysis were conducted for the purpose of gaining comprehension of the internal representation developed. We applied Principle Component Analysis (PCA) to the estimated posterior latent states corresponding to the command voice input, incorporating the tasks and colors stated in each command. Figure 5 presents evidence that robots with all curiosity developed a a compositional and generalizable understanding of these goals. At the midpoint of training, the latent representations begin to show consistent grouping of tasks and colors. For example, the tasks “watch,” “be near,” and “push forward” are tightly grouped, suggesting that these tasks are interpreted as similar. In contrast, the task “push right” appears heavily separated from other tasks, and the tasks “touch the top” and “push left” also appear as distinct categories. After training, these clusters of tasks are more compact and separated. Within these clusters, there is loose sub-structuring by color: green and yellow tend to be on the left side of the cluster, while blue and magenta are on the right side of the cluster. Therefore, it can be said that each cluster represents a distinct linguistic concept while exhibiting relationship with others since the learning of visuo-proprioceptive-motor also contributes to this structuring. In contrast, Figure 6 shows the same type of PCA results for a robot with no curiosity. “Be near,” “push forward,” and “push right” are heavily entangled with each other; coincidentally, in Figure 3, these are the three tasks which these robots perform worst. Thus supports the idea that curiosity aids task disentanglement and compositional learning.
Figure 5: Compositional understanding in robots with all curiosity. PCA of latent variables corresponding to the command voice. Left: halfway through training. Right: after training. Clusters with substructures emerged early and became more refined over time.
Figure 6: Compositional understanding in robots with no curiosity. PCA applied to latent representations of command voice inputs after training. Tasks with poor performance remain entangled.
In the current model, the robot’s knowledge in the environment should develop richer along the course of exploratory learning. For confirmation of this idea, we examined the capability of the robots in generating mental plans for achieving goals without accessing the sensory inputs except the initial step for an episode trial as compared between the half-trained case and the fully trained case. A robot’s mental planning can be visualized by allowing it to receive real sensory observation only at the initial step, after which the robot must rely entirely on its own internal predictions. In this setting, the robot views its predicted sensory observations as if they are true inputs. This process may be likened to a dreamlike state or hallucinatory simulation, in which the robot mentally simulates future events based on its internal model of the world. Figure 7 (a) illustrates a simulation example for the fully trained case. The robot was commanded to touch the top of the yellow pillar. In that figure, the first row shows the ground truth environment from a view behind the robot’s shoulder. The second row shows what the robot would truly observe if it were not in this planning setting. The third row shows the robot’s look ahead predictions for visual observations, which it will interpret as if they are real. It can be seen that even with only the first step sensory observation, the robot could generate mostly accurate future look-ahead prediction for sensation as well as motor command. These predictions are sufficiently accurate for the robot to maintain an internal conceptualization of the environment and complete its command in the case of the end of the developmental learning. Figure 7 (b) illustrates the same robot after only half of its training in the same scenario. In this case, the robot’s predictions are inaccurate, causing it to wander and view an object which does not actually exist.
Figure 7: A mentally planning robot commanded to touch the top of the yellow pillar. The first row displays the ground truth from a view over the robot’s right shoulder. The second row displays the vision which the robot would have experienced if it were not only mentally planning; the robot observes only the first genuine sensory observation. Subsequent inputs are the robot’s own predictions, which it treats as real, displayed in the third row. (A) Despite the limitation, the robot successfully completes the commanded action and maintains a consistent internal model of its environment. (B) The same robot in the same scenario, but after only half of its training, is unable to make correct predictions and therefore cannot complete the command action.

### v3

# Curiosity-Driven Development of Action and Language in Robots Through Self-Exploration

@@ Abstract
Abstract
Human infants acquire language and action gradually through development, achieving strong generalization from minimal experience, whereas large language models require exposure to billions of training tokens. What mechanisms underlie such efficient developmental learning in humans? This study investigates this question through robot simulation experiments in which agents learn to perform actions associated with imperative sentences (e.g., push red cube) via curiosity-driven self-exploration. Our approach integrates the active inference framework with reinforcement learning, enabling intrinsically motivated developmental learning. The simulations reveal several key findings: i) Generalization improves markedly as the scale of compositional elements increases. ii) Curiosity combined with motor noise yields substantially better learning than exploration without curiosity. iii) Rote pairing of sentences and actions precedes the emergence of compositional generalization. iv) Simpler, prerequisite-like actions develop earlier than more complex actions that depend on them. v) When exception-handling rules were introduced—where certain imperative sentences required executing inconsistent actions—the robots successfully acquired these exceptions through exploration and displayed a U-shaped performance curve characteristic of representational redescription in child language learning. Together, these results suggest that curiosity-driven exploration and active inference provide a powerful account of how intrinsic motivation and hierarchical sensorimotor learning can jointly support scalable compositional generalization and exception handling in both humans and artificial agents.
@@ Introduction
A central question in both cognitive science and artificial intelligence is how humans and artificial systems can acquire competencies for language and action developmentally, despite having access to only limited learning experiences. This question is exemplified in human infants, who achieve remarkable generalization with sparse input. This is a stark contrast to large-scale models which rely on massive training corpora to reach similar capabilities. This raises the issue of what mechanisms enable such efficient developmental learning.
@@
In linguistic terms, compositionality refers to the ability to construct novel configurations by systematically combining elements such as verbs, adjectives, and nouns. Generalization enables infants to apply learned components flexibly, allowing for the production and interpretation of utterances that have not been directly encountered previously. Although the number of possible compositions grows multiplicatively with the vocabulary size (i.e., number of verbs $\times$ number of adjectives $\times$ number of nouns), infants achieve generalization after experiencing only a small subset of learning examples. This suggests that the effective sample complexity could be proportional to the sum of elements rather than their product. This phenomenon is closely related to the “poverty of the stimulus” problem articulated by Chomsky (?), which asks how learners generalize so effectively given severely sparse input.
@@
To investigate these mechanisms, one promising approach is to reconstruct developmental learning processes in machines and robots. The field of developmental robotics has long pursued this line of research, aiming to replicate human-like learning trajectories in embodied systems (?, ?, ?, ?). However, relatively few studies have focused on development of language and motor control under conditions of stimulus poverty. Existing work has primarily examined associative mappings between linguistic input and motor commands in one-shot or supervised batch learning schemes (?, ?, ?, ?). These approaches neglect the self-directed, developmental context of infant learning.
@@
In this study, we propose a self-exploratory learning framework of robots in which reinforcement learning is incorporated with the active inference framework (?, ?, ?), enabling curiosity-driven exploration. Our approach to integrate reinforcement learning with active inference was originally inspired by the work of Kawahara et al. (?). In our model, originally introduced in (?), motor commands are reinforced by two intrinsic rewards: curiosity (seeking unpredictable sensory consequences) and motor entropy (seeking random movements). Motor commands are also reinforced by extrinsic rewards for successfully achieving goal verbs specified by given imperative sentences. Importantly, our previous experiments in maze navigation demonstrated that the combination of curiosity and motor entropy is crucial for enhancing self-exploration, as agents achieved significantly improved exploratory behaviors under this dual-intrinsic reward scheme. Our approach aligns with broader research on self-exploration in machine learning, in which agents are intrinsically rewarded for taking motor commands that increase unpredictability or information gain (?, ?).
@@
A simulated mobile robot equipped with a manipulator arm, vision sensor, and distributed tactile sensors learns to generate motor movements in response to imperative sentences presented during each trial. These sentences are systematically composed of verbs, adjectives, and nouns, enabling evaluation of generalization performance under different levels of compositional complexity.
Beyond these primary developmental processes, we further explore the capacity for exception handling—a hallmark of flexible cognition. In human development, exceptions such as irregular verbs or inconsistent mappings often produce non-monotonic, U-shaped learning trajectories: children first apply a correct form, then overgeneralize it (producing errors), and finally recover the correct rule. This pattern has been widely interpreted as evidence of internal representational reorganization or representational redescription (?). Computationally, such U-shaped performance has been demonstrated in models of language acquisition and rule learning (?, ?, ?, ?, ?). Developmentally, these phenomena reflect the tension between rote memorization, generalization, and the later refinement of exception rules.
To examine whether similar mechanisms can emerge in artificial developmental systems, we hypothesize that robots trained through curiosity-driven active inference can also acquire exception-handling rules through exploratory learning. Specifically, we predict that when presented with a small number of abnormal linguistic rules—commands requiring reversed or inconsistent actions—the robot will exhibit a U-shaped performance curve: initial correct learning, followed by a decline due to overgeneralization, and eventual recovery as exception-specific representations are internalized. This pattern would indicate that the system develops not merely through associative reinforcement, but through representational reorganization analogous to cognitive processes observed in human infants.
@@
This study therefore tested the following hypotheses through simulation experiments: H1: Generalization performance improves as the scale of compositionality in the task increases. H2: Curiosity combined with motor entropy enhances the performance of developmental learning. H3: In the early phase, actions are generated only for exactly learned imperative sentences, but in later phases, the system generalizes to novel, unlearned compositions. H4: Primitive actions are acquired earlier, followed by more complex, prerequisite-dependent actions. H5: Exception-handling rules can be acquired through exploratory learning, exhibiting U-shaped developmental performance similar to that observed in human cognition.
@@ Results
We created a robot like a truck crane in a physics simulator along with a set of objects with 5 different shapes each of which can be with 6 different colors (see Fig. 1).
Fig. 1: The simulated robot and a set of objects to act on. (A) The robot has two wheels and an arm with two joints. The design is similar to a truck crane. (B) Left to right: a red pillar, a green pole, a blue dumbbell, a cyan cone, and a magenta hourglass. The color yellow is not pictured here.
@@
For each trial episode, a task goal was given in terms of an imperative sentence composed with verb, adjective, and noun. Possible words used for them are shown in Table 1.
@@
Verb Adjective Noun
@@
Table 1: English words. The English words used for imperative sentences specifying goals.
@@
At each step, the robot receives visual sensation, proprioception for the arm, tactile sensation , and two types of voices: the command voice and the tutor-feedback voice. The command voice takes the format of the imperative sentence described previously, and it comes every step continuously from the beginning. On the other hand, the feedback voice arrives whenever the robot achieves one of possible goals even if the achieved goal is not the imperative sentence told by the command voice, and it informs which goal has been achieved actually in the same format with the command voice. This potentially enhances the forward model to learn about own action as associated with linguistic representation. Finally, when the goal specified by the command voice is achieved, a reward is provided. Each trial episode ran for 30 steps, or terminated when the specified goal is achieved.
@@
The experiment examined effects of different levels of curiosity to the developmental learning processes using the basic setup. In the basic setup, full compositions of words (Table ) were used to generate the imperative sentences. However, the training was conducted using only 60 imperative sentences (33 percentage) out of 180 possible sentences. 120 untrained sentences were used for generalization test. For ten robots with different random seeds, the complete developmental learning process was iterated for 60000 epochs. The generalization test with unlearned imperative sentences was conducted for every 50 epochs.
@@
Fig. 2 shows the development of the generalization test performances in terms of success rate for goals specified by unlearned imperative sentences which are plotted for different action categories with different levels of curiosity. Shades areas represent 99% confidence intervals.
@@
Fig. 2: Rolling success-rates for unlearned goals. Compares agents with different levels of curiosity.
@@
The plots shows that the performance was improved significantly as the curiosity level was increased. Especially, the case of all actions with the all curiosity level shows the average success rate for unlearned goals reached a quite high value of 85 percentage even though the learning was conducted only for 33 percentage of all possible compositions. Video Experiment 2 shows examples of the behaviors of robots with the all curiosity level. In the intermediate phase of development, the robot acts with play-like behavior. In the final phase of development, the robot quickly and accurately achieves its goals. It can be also seen in Fig. 2 that some action categories developed faster than other action categories. Especially in the all curiosity case, “watch” developed the fastest, “be near” did as the second, “touch the top” as the third, and “push left” “push right”, and “push forward” developed much later. This implies that simpler prerequisite-type actions develop earlier, and more complex actions requiring those prerequisite actions develop later. Actually, an action of watching an object should be prerequisite for all other object-targeted actions including an action of moving near by an object, which should be prerequisite again for actions of directly manipulating an object like pushing left/right an object or touching the top of it. Our observation accords with this.
@@
Next, Fig. 6 (A) shows the success rate comparison between learned and unlearned goals under the all curiosity condition for each action category. These plots show that the test performance for learned goals developed significantly faster than the case for the unlearned goals. This indicates that actions are generated only for exactly learned compositional imperative sentences in the early phase, but the system generalizes to novel, unlearned ones in the later phase.
Further analysis
Some analysis were conducted for the purpose of gaining comprehension of the internal representation developed. We applied Principle Component Analysis (PCA) to the estimated posterior latent states corresponding to the command voice input, incorporating the verbs and adjectives stated in each command. Fig. 3 presents evidence that robots with all curiosity developed a compositional and generalizable understanding of these goals. At the midpoint of training, the latent representations begin to show consistent grouping of verbs and adjectives. For example, the verbs “watch,” “be near,” and “push forward” are tightly grouped, suggesting that these verbs are interpreted as similar. In contrast, the verb “push right” appears heavily separated from other verbs, and the verbs “touch the top” and “push left” also appear as distinct categories. After training, these clusters of verbs are more compact and separated. Within these clusters, there is loose sub-structuring by color: green and yellow tend to be on the left side of the cluster, while blue and magenta are on the right side of the cluster. Therefore, it can be said that each cluster represents a distinct linguistic concept while exhibiting relationship with others since the learning of visuo-proprioceptive-motor also contributes to this structuring. In contrast, Fig. 4 shows the same type of PCA results for a robot with no curiosity. “Be near,” “push forward,” and “push right” are heavily entangled with each other. Thus supports the idea that curiosity aids verb disentanglement and compositional learning.
Fig. 3: PCA for language latent variables in the case of all curiosity. PCA of latent variables corresponding to the command voice. (A) Halfway through development. (B) After complete development. Clusters with substructures emerged early and became more refined over time.
Fig. 4: PCA for language latent variables in the case of no curiosity. PCA applied to latent representations of command voice inputs after complete development.
In the current model, the robot’s knowledge of the environment should develop richer along the course of exploratory learning. For confirmation of this idea, we examined the capability of the robots in generating mental plans for achieving goals without accessing the sensory inputs except the initial step for an episode trial as compared between the half-trained case and the fully trained case. A robot’s mental planning can be visualized by allowing it to receive real sensory observation only at the initial step, after which the robot must rely entirely on its own internal predictions. In this setting, the robot views its predicted sensory observations as if they are true inputs. This process may be likened to a dreamlike state or hallucinatory simulation, in which the robot mentally simulates future events based on its internal model of the world. Fig. 5 (A) illustrates a simulation example for the fully trained case. The robot was commanded to touch the top of the yellow pillar. In that figure, the first row shows the ground truth environment from a view behind the robot’s shoulder. The second row shows what the robot would truly observe if it were not in this planning setting. The third row shows the robot’s look ahead predictions for visual observations, which it will interpret as if they are real. It can be seen that even with only the first step sensory observation, the robot could generate mostly accurate future look-ahead prediction for sensation as well as motor command. These predictions are sufficiently accurate for the robot to maintain an internal conceptualization of the environment and complete its command in the case of the end of the developmental learning. Fig. 5 (B) illustrates the same robot after only half of its training in the same scenario. In this case, the robot’s predictions are inaccurate, causing it to wander and view an object which does not actually exist.
Fig. 5: Mental plans generated by the robot. This robot is commanded to touch the top of the yellow pillar. The first row displays the ground truth from a view over the robot’s right shoulder. The second row displays the visual sequence of the ground truth. The third row shows the visual sequence of mental planning. (A) The case for the end of complete developmental learning, and (B) for the case of halfway developed.
@@
The experimental results are shown in Fig. 6. Shades areas represent 99% confidence intervals. It can be seen that although the learned goal test cases show equally high performance for all scales of compositionality, the generalization test for unlearned goal case shows that the success rate in the final trial decreases significantly (85 percentage to 25 percentage) as the compositionality scale decreases. This indicates that the generalization performance severely depends on the scale of compositionality in learning examples.
Fig. 6: Rolling success-rates for learned and unlearned goals with different compositionality scales. (A) Agents trained with all six verbs, all six adjectives, and all five nouns. (B) Agents trained with five verbs, five adjectives, and three nouns. (C) Agents trained with four verbs, four adjectives, and three nouns.
Name Verbs Adjectives Nouns
Largest Vocabulary
Watch
Be Near
Touch the Top
Push Forward
Push Left
Push Right
Red
Green
Blue
Cyan
Magenta
Yellow
Pillar
Pole
Dumbbell
Cone
Hourglass
Reduced Vocabulary
Watch
Be Near
Push Forward
Push Left
Push Right
Red
Green
Blue
Cyan
Magenta
Pillar
Pole
Dumbbell
Cone
Smallest Vocabulary
Watch
Push Forward
Push Left
Push Right
Red
Green
Blue
Cyan
Pillar
Pole
Dumbbell
Table 2: Verbs, adjectives, and nouns which are used for training agents in three different ways.
Exception rule handling: Experiment 3
This experiment examined how robots can acquire exception handling rules through developmental learning. While most command–action mappings were preserved, the commands “watch magenta pillar” and “be near green pole” were swapped: success required performing the other goal, not the one commanded. These mismatches required the robot to override its learned generalized knowledge.
This simulation experiment was conducted using the same model parameters with the ones used in the previous experiments with 10 robots with 60,000 epochs of developmental trials. At the end of development the average success rate among 10 robots was 84 percentage for the learned goals, 76 percentage for unlearned goals, and 50 percentage for the exception handling cases. The average success rate for the exception handling cases is not so high most likely due to over-generalization.
Panel A in Fig. 7 shows the rolling success-rate for achieving the exception goals for each of 10 individual robots while Panel B shows the success rate for the same goals but without applying the exception handling rules. The final success rate for the exception handling case is diverse ranging from 15 percentage to 90 percentage as can be seen in Panel A. It was also discovered that 7 out of 10 robots trained with these exceptions exhibit characteristic U-shaped curves: early success, followed by a drop, and eventual recovery with higher success rate than the earlier one. In contrast, monotonic increase of success rate can be seen in all 10 individuals in the case of learning without the exception handling rules. Statistical comparisons (detailed in the Supplementary Text) confirm that U-shaped patterns are significantly more prevalent in the exception condition than in the control ($p=.0025$).
In Fig. 8, PCA illustrates how the internal representations of of goals in the seventh robot in Fig. 7 A evolve in a manner consistent with the U-shaped success rates. In panel A, early in training, goal embeddings are muddled without clear structure, reflecting a learning phase with minimal generalization. In panel B, midway in training, the exception command “watch magenta pillar” is embedded near other “watch” goals, while “be near green pole” clusters with other “be near” goals, despite these associations being incorrect. These observation indicates that overgeneralization has occurred. In panel C, late in training, “watch magenta pillar” is now embedded near “be near” goals, and “be near green pole” near “watch” goals, indicating that the robot has correctly handled these exceptions as swapped pairs. Here, it can be said that the representational redescription took place in the course of developmental learning of exception handling rules.
Fig. 7: Comparison of the performance curve with and without applying the exception handling rules. (A) The development of success rate in achieving two goals which are swapped as exceptions. Red vertical lines depict the peaks and valleys of the learning, as defined in supplementary text. (B) The development of success rate in achieving the same two goals without applying the exception handling rules.
Fig. 8: PCA for language latent variables for an individual developed with the exception handling rules. (A) Plot at 20000 epochs, (B) plot at 32500 epochs, and (C) plot at 57500 epochs wherein the circled ”N” and circled ”W” denote the sentences applied with the exception handling